# Cross-positive linear maps, positive polynomials and sums of squares

### v1

@@ Abstract.
A linear map $\Phi$ between matrix spaces is called cross-positive if it is positive on orthogonal pairs $(U,V)$ of positive semidefinite matrices in the sense that $\langle U,V\rangle:=\operatorname{tr}(UV)=0$ implies $\langle\Phi(U),V\rangle\geq 0$, and is completely cross-positive if all its ampliations $I_{n}\otimes\Phi$ are cross-positive. (Completely) cross-positive maps arise in the theory of operator semigroups, where they are sometimes called exponentially-positive maps, and are also important in the theory of affine processes on symmetric cones in mathematical finance.
@@
${}^{1}$upported by the Slovenian Research Agency program P1-0222 and grants J1-50002, J1-8132, J1-2453, N1-0217, J1-3004.
${}^{2}$upported by the Slovenian Research Agency program P1-0222 and grants J1-8132, N1-0217, J1-2453, J1-3004.
${}^{3}$upported by the Slovenian Research Agency grants J1-50002, J1-2453, J1-8132, J1-3004 and P1-0288.
@@ 1. Introduction
We use $M_{n}(\mathbb{R})_{\textrm{sym}}$ to denote the set of all real symmetric $n\times n$ matrices, and $M_{n}(\mathbb{R})_{\succeq 0}$ for the set of all positive semidefinite (symmetric) matrices. We let $I_{n}$ (resp. $0_{n}$) stand for the $n\times n$ identity (resp. zero) matrix. A linear map $\varphi:M_{n}(\mathbb{R})\to M_{n}(\mathbb{R})$ is called positive if it maps positive semidefinite matrices into positive semidefinite matrices, and is completely positive if all its ampliations $I_{m}\otimes\varphi$ ($m\in\mathbb{N}$) are positive. Consider a linear map $A:M_{n}(\mathbb{R})\to M_{n}(\mathbb{R})$. For each $t\in\mathbb{R}$ the linear map $\exp(tA):M_{n}(\mathbb{R})\to M_{n}(\mathbb{R})$ is defined by $\exp(tA)=\sum_{i=0}^{\infty}\frac{1}{i!}(tA)^{i}$. The operator valued function $t\mapsto\exp(tA)$ is the solution of the differential equation $\dot{X}(t)=AX(t)$, which makes it important in analysis and applications to physics [Lin76], math finance [CFMT11, CK-RMT16], etc. The well-known formula $\exp((s+t)A)=\exp(sA)\circ\exp(tA)$ implies that the set $\{\exp(tA)\colon t\geq 0\}$ is a (one-parameter) semigroup under composition. The linear map $A$ is called the generator of the one-parameter semigroup. We call $A$ exponentially-positive, resp. completely exponentially-positive, if $\exp(tA)$ is a positive, resp. completely positive map for all $t\geq 0$. In such a case the semigroup $\{\exp(tA)\colon t\geq 0\}$ is called a positive, resp. completely positive one-parameter semigroup. Note that positivity of linear maps and one-parameter semigroups of them is studied more generally over any ordered vector spaces, in finite and infinite dimensions, and for bounded and unbounded linear operators. We refer the reader to [EH-O79, Nag86] for detailed studies.
The exponential positivity property can be rephrased in a more traditional matrix theory form as follows:
Theorem 1.1 (Schneider, Vidyasagar [SV70, Theorem 3]).
A linear map $A:M_{n}(\mathbb{R})\to M_{n}(\mathbb{R})$ is exponentially-positive iff
(1.1) $\forall U,V\in M_{n}(\mathbb{R})_{\succeq 0}:\langle U,V\rangle=0\quad\Rightarrow\quad\langle A(U),V\rangle\geq 0.$
[…]
A linear map satisfying (1.1) is cross-positive. Similarly, we call $A$ completely cross-positive if
@@
Thus, by Theorem 1.1 and the equality $\exp(t(I_{k}\otimes A))=I_{k}\otimes\exp(tA)$, $A$ is completely cross-positive iff $A$ is completely exponentially-positive. Indeed, by definition, $A$ is completely cross-positive if and only if $I_{k}\otimes A$ is cross-positive for each $k\in\mathbb{N}$. By Theorem 1.1 this holds if and only if $I_{k}\otimes A$ is exponentially-positive for each $k\in\mathbb{N}$, i.e., if and only if
$\forall k\in\mathbb{N},\forall t\geq 0,\forall X\succeq 0$ ::: $\exp(t(I_{k}\otimes A))(X)=\sum_{i=0}^{\infty}\frac{1}{i!}t^{i}(I_{k}\otimes A)^{i}(X)$
@@
However, this is equivalent to complete positivity of $\exp(tA)$ for each $t\geq 0$, i.e., to complete exponential positivity of $A$.
In [KOŠT15] the authors construct, for the first time, a proper cross-positive map, that is, a cross-positive map that is not completely cross-positive. Such maps and the associated one-parameter semigroups of endomorphisms of a symmetric cone are an important ingredient in the theory of affine processes on symmetric cones. Affine processes play a major role in math finance [CK-RMT16]; they are simple enough to be tractable from the point of view of theory and numerics, while at the same time sufficiently flexible from a modeling point of view. Affine processes on the cone of real positive semidefinite matrices were classified in [CFMT11, Theorem 2.4], see also [CK-RMT16, Theorem 2.19] for the classification of affine processes on all symmetric cones. According to the classification, the linear drift of an affine process is given by a cross-positive map. The cross-positive map defining the drift is unique only modulo an integral with respect to a measure that describes jumps of the affine process. The operator defined by the integral is completely positive, so a drift defined by a cross-positive, but not completely cross-positive map cannot be removed by a change of measure. See [CK-RMT16] or [KOŠT15, Section 6] for more details.
In this paper we investigate and quantify the gap between cross-positive maps and completely cross-positive maps, and provide an algorithm for providing further examples of proper cross-positive maps. In addition to matrix analysis our main tools include real algebraic geometry [BCR98], convexity [Scn93, BB05] and harmonic analysis [Duo87].
1.1. Main results and readers’ guide
In the preliminary Section 2 we translate the properties of linear maps $A:M_{n}(\mathbb{R})\to M_{n}(\mathbb{R})$ to properties of biquadratic forms
(1.3) $p_{A}:={\tt y}^{T}A({\tt x}{\tt x}^{T}){\tt y}\in\mathbb{R}[{\tt x},{\tt y}],$
where ${\tt x}=({\tt x}_{1},\ldots,{\tt x}_{n}),$ ${\tt y}=({\tt y}_{1},\ldots,{\tt y}_{n})$ are tuples of commuting indeterminates. Then we explain that completely cross-positive maps are much tamer and easier to handle than cross-positive maps, resembling the well-known relationship between positive and completely positive maps [Cho75, Arv09, KMŠZ19, tHvdM21].
The main contribution of this article is three-fold. First, we quantify the gap between cross-positive and completely cross-positive maps. Roughly speaking, very few cross-positive maps are completely cross-positive. More precisely, as shown in Corollary 4.10, the probability $p_{n}$ that a cross-positive map $M_{n}(\mathbb{R})_{\textrm{sym}}\to M_{n}(\mathbb{R})_{\textrm{sym}}$ is completely cross-positive, is $p_{n}<(Cn)^{-\frac{1}{2}\binom{n+1}{2}^{2}}$ for an absolute constant $C$, so $\lim\limits_{n\to\infty}p_{n}=0.$ Our proof (in Appendices A and B) roughly follows Blekherman’s outline in his papers characterizing the gap between positive and sum of squares polynomials [Ble06, BB05]. A key new ingredient in the proof is a dimension-independent reverse Hölder inequality for bilinear biforms given in Section 4.3.
Second, Section 3 considers the smallest nontrivial case, that is, the case of cross-positive maps $A:M_{3}(\mathbb{R})_{\textrm{sym}}\to M_{3}(\mathbb{R})_{\textrm{sym}}$. We give real algebraic geometry inspired certificates (Nichtnegativstellensätze) for $A$ to be cross-positive; see Theorem 3.3 for the case $A$ satisfies some mild nonsingularity-type assumption, and Corollary 3.10 for the singular case.
Finally, in Section 5, as a side product of our analysis we provide a randomized polynomial-time algorithm based on semidefinite programming [WSV00] for producing proper cross-positive maps.
2. Preliminaries
Let $n\geq 2$. To each linear map $A:M_{n}(\mathbb{R})_{\textrm{sym}}\to M_{n}(\mathbb{R})_{\textrm{sym}}$ we assign the biquadratic form $p_{A}\in\mathbb{R}[{\tt x},{\tt y}]$ as in (1.3). Let $I\subseteq\mathbb{R}[{\tt x},{\tt y}]$ be the ideal generated by ${\tt y}^{T}{\tt x}=\sum_{i=1}^{n}{\tt x}_{i}{\tt y}_{i}$, and let $V(I)$ be the corresponding real variety
@@ 2. Preliminaries
The variety $V(I)$ is an irreducible hypersurface for $n\geq 2$ and the defining polynomial ${\tt y}^{T}{\tt x}$ changes sign on $\mathbb{R}^{2n}$. Hence the ideal $I$ is real radical [BCR98, Theorem 4.5.1]. Thus $I$ is the vanishing ideal of $V(I)$, i.e., a polynomial $p\in\mathbb{R}[{\tt x},{\tt y}]$ vanishes on $V(I)$ iff $p\in I$.
The following is a special case of [KOŠT15, Corollary 15] and [TG13, Theorem 2], but can also be established by a straight-forward calculation.
@@
[…]
Remark 2.2.
A sum of a positive map and a map of the form (2.1) is clearly cross-positive. The converse is true up to closure, see [SV70, Lemma 6 and Theorem 2]. It has been conjectured for some time that each cross-positive map actually is a sum of a positive map and a map of the form (2.1) (see [Damm04, Section 4] or [CFMT11, p.409]), but a counterexample was found in [KOŠT15]. Such counterexamples were called exotic cross-positive maps in [KOŠT15]. On the other hand, an analogous counterexample does not exist for completely cross-positive maps, see Proposition 2.5 below.
@@ 2.1. Parametrization of $V(I)$
(i)$\Leftarrow$(ii) Assume $p_{A}$ is nonnegative on $V(I).$ Given $U,V\in M_{n}(\mathbb{R})_{\succeq 0}$ with $\langle U,V\rangle=0$, write $U=\sum u_{i}u_{i}^{T}$ and $V=\sum v_{i}v_{i}^{T}$. As the scalar product of two positive semidefinite matrices is nonnegative, we deduce $\langle u_{i},v_{j}\rangle=0$ for all $i,j$. The assumption now implies $p_{A}(u_{i},v_{j})\geq 0$. Then
@@
[…]
@@
    $A$ is completely cross-positive;
@@
Let ${\underline{r}}:=(r_{1},\ldots,r_{n})\in\mathbb{Z}^{n}_{+}$ stand for a tuple of non-negative integers, ${\tt x}^{\underline{r}}$ for the monomial ${\tt x}_{1}^{r_{1}}\cdots{\tt x}_{n}^{r_{n}}$ and $\operatorname{conv}(E)\subseteq\mathbb{R}^{n}$ for the convex hull of the set $E\subseteq\mathbb{R}^{n}$. Recall that the Newton polytope $N(p)$ of a polynomial $p({\tt x})=\sum_{{\underline{r}}}c_{{\underline{r}}}{\tt x}^{{\underline{r}}}$, where $c_{\underline{r}}\in\mathbb{R}\setminus\{0\}$, is the convex hull of the exponent vectors of the monomials appearing nontrivially in $p$, i.e.,
@@
(i)$\Leftrightarrow$(ii) It is well-known that a linear map on $M_{n}(\mathbb{R})_{\textrm{sym}}$ is completely cross-positive iff it is a sum of a completely positive map and a map of the form (2.1), see [Lin76, Theorem 3] for a nice proof of this result. Using Lemma 2.1 and [KMŠZ19, Proposition 3.1] we can then see that $A$ is completely cross-positive iff $p_{A}=\sum_{i=1}^{k}p_{i}^{2}+q$ for some bilinear forms $p_{i}$ and some biquadratic form $q\in I$, i.e., $p_{A}$ is a sum of squares modulo $I$ by Lemma 2.3.
(ii)$\Rightarrow$(iii) is obvious, so we prove (iii)$\Rightarrow$(ii). First note that multi-homogeneity of $q_{A}$ implies that $q_{A}$ is a sum of squares of biforms that are quadratic in ${\tt x}$ and linear in $\alpha$. Write
@@
where $i,\ell=2,\ldots,n-1$, $j,k=1,\ldots,n$. By [Rez78, Theorem 1], we have the inclusions $N(q^{(\ell)})\subseteq\frac{1}{2}N(q_{A})$ of Newton polytopes, which implies that each $q^{(\ell)}$ is a linear combination of the monomials
@@
Proof of Claim. We have to consider how each of the monomials in (2.6) can appear in $q^{(\ell)}({\tt x},\alpha)$. For $j=1,\ldots,n$, the monomials
@@
If any of the monomials from (2.7) occurs in $q^{(\ell)}({\tt x},\alpha)$, then it also occurs in the polynomial $q^{(\ell)}(\widehat{\tt x}_{i-1},\widehat{\alpha}_{i-1})$ with the same coefficient. By definition we have that
@@
Hence, each $q^{(\ell)}(\widehat{{\tt x}}_{i-1},\widehat{\alpha}_{i-1})$ vanishes on $V(\alpha_{i}{\tt x}_{i+1}-\alpha_{i-1}{\tt x}_{i-1})$. Since $\alpha_{i}{\tt x}_{i+1}-\alpha_{i-1}{\tt x}_{i-1}$ is irreducible in $\mathbb{R}[{\tt x},\alpha]$ and it changes sign on $\mathbb{R}^{2n-1}$, it follows by [BCR98, Theorem 4.5.1] that
@@
where $p_{i-1}$ is a linear form in $\widehat{{\tt x}}_{i-1}$. Now (2.8) implies that the monomials from (2.7) can appear non-trivially in $q^{(\ell)}(\widehat{{\tt x}}_{i-1},\widehat{\alpha}_{i-1})$ only from the scalar multiple of the term
@@
Using the Claim and (2.5) it follows that $p_{A}({\tt x},{\tt y})$ agrees on a dense subset of $V(I)$ and by continuity on the whole $V(I)$ with an sos polynomial, which we denote by $r({\tt x},{\tt y})$. Since $p_{A}-r$ vanishes on $V(I)$, the polynomial ${\tt y}^{T}{\tt x}$ is irreducible in $\mathbb{R}[{\tt x},{\tt y}]$ and its sign changes on $\mathbb{R}^{2n-1}$, it follows that $p_{A}-r\in I$ by [BCR98, Theorem 4.5.1]. ∎

### v2

@@ Abstract.
A $\ast$-linear map $\Phi$ between matrix spaces is cross-positive if it is positive on orthogonal pairs $(U,V)$ of positive semidefinite matrices in the sense that $\langle U,V\rangle:=\operatorname{tr}(UV)=0$ implies $\langle\Phi(U),V\rangle\geq 0$, and is completely cross-positive if all its ampliations $I_{n}\otimes\Phi$ are cross-positive. (Completely) cross-positive maps arise in the theory of operator semigroups, where they are sometimes called exponentially-positive maps, and are also important in the theory of affine processes on symmetric cones in mathematical finance.
@@
1Supported by the Slovenian Research Agency program P1-0222 and grants J1-50002, J1-8132, J1-2453, N1-0217, J1-3004, J1-60011.
2Supported by the Slovenian Research Agency program P1-0222 and grants J1-8132, N1-0217, J1-2453, J1-3004, J1-60011.
3Supported by the Slovenian Research Agency grants J1-50002, J1-2453, J1-8132, J1-3004, J1-60011 and P1-0288.
[…]
@@ 1. Introduction
Let $M_{n}(\mathbb{R})$ be the vector space of $n\times n$ real matrices equipped with the involution $T$ which is the usual transposition of matrices. We use $M_{n}(\mathbb{R})_{\succeq 0}$ to denote the set of all positive semidefinite (symmetric) matrices. We let $I_{n}$ (resp. $0_{n}$) stand for the $n\times n$ identity (resp. zero) matrix. A linear map $A:M_{n}(\mathbb{R})\to M_{n}(\mathbb{R})$ is $\ast$-linear if $A(U^{T})=A(U)^{T}$ for all $U\in M_{n}(\mathbb{R})$. A $\ast$-linear map $A$ is positive if it maps positive semidefinite matrices into positive semidefinite matrices, and is completely positive if the ampliations
$I_{k}\otimes A:M_{k}(\mathbb{R})\otimes M_{n}(\mathbb{R})\to M_{k}(\mathbb{R})\otimes M_{n}(\mathbb{R}),\quad U\otimes V\mapsto U\otimes A(V)$
are positive for every $k\in\mathbb{N}$. Here $\otimes$ stands for the Kronecker tensor product of matrices. Relaxing positivity of $A$ to the condition
(1.1) $\forall U,V\in M_{n}(\mathbb{R})_{\succeq 0}:\langle U,V\rangle=0\quad\Rightarrow\quad\langle A(U),V\rangle\geq 0,$
where $\langle\textvisiblespace,\textvisiblespace\rangle$ denotes the standard scalar product on $M_{n}(\mathbb{R})$, i.e., $\langle B,C\rangle:=\operatorname{tr}(C^{T}B)$, gives a definition of a cross-positivity of $A$ in which case $A$ is cross-positive. Similarly, we call $A$ completely cross-positive if
@@
In [23] the authors construct, for the first time, a proper cross-positive map $A$, that is, a cross-positive map that is not completely cross-positive. Such maps and the associated one-parameter semigroups (under composition) $\{\exp(tA)\colon t\geq 0\}$ of endomorphisms of a symmetric cone are an important ingredient in the theory of affine processes on symmetric cones. In the semigroup theory cross-positive (resp. completely cross-positive) maps are known as exponentially-positive (resp. completely exponentially-positive) maps (see Section 2.1 for details). Affine processes play a major role in math finance [14]; they are simple enough to be tractable from the point of view of theory and numerics, while at the same time sufficiently flexible from a modeling point of view. Affine processes on the cone of real positive semidefinite matrices were classified in [13, Theorem 2.4], see also [14, Theorem 2.19] for the classification of affine processes on all symmetric cones. According to the classification, the linear drift of an affine process is given by a cross-positive map. The cross-positive map defining the drift is unique only modulo an integral with respect to a measure that describes jumps of the affine process. The operator defined by the integral is completely positive, so a drift defined by a cross-positive, but not completely, cross-positive map cannot be removed by a change of measure. See [14] or [23, Section 6] for more details.
In this paper we investigate and quantify the gap between cross-positive maps and completely cross-positive maps, and provide an algorithm for providing further examples of proper cross-positive maps. In addition to matrix analysis our main tools include real algebraic geometry [8], convexity [37, 3] and harmonic analysis [16].
1.1. Main results and readers’ guide
In the preliminary Section 2 we translate the properties of $\ast$-linear maps $A:M_{n}(\mathbb{R})\to M_{n}(\mathbb{R})$ to properties of biquadratic forms
(1.3) $p_{A}:={\tt y}^{T}A({\tt x}{\tt x}^{T}){\tt y}\in\mathbb{R}[{\tt x},{\tt y}],$
where ${\tt x}=({\tt x}_{1},\ldots,{\tt x}_{n}),$ ${\tt y}=({\tt y}_{1},\ldots,{\tt y}_{n})$ are tuples of commuting indeterminates. Then we explain that completely cross-positive maps are much tamer and easier to handle than cross-positive maps, resembling the well-known relationship between positive and completely positive maps [12, 1, 22, 40].
The main contribution of this article is three-fold. First, we quantify the gap between cross-positive and completely cross-positive maps. Roughly speaking, very few cross-positive maps are completely cross-positive. More precisely, as shown in Corollary 4.8, the probability $p_{n}$ that a cross-positive map $M_{n}(\mathbb{R})\to M_{n}(\mathbb{R})$ is completely cross-positive, is less than $(Cn)^{-\frac{1}{2}\binom{n+1}{2}^{2}}$ for an absolute constant $C$, so $\lim\limits_{n\to\infty}p_{n}=0.$ Our proof roughly follows Blekherman’s outline in his papers characterizing the gap between positive and sum of squares polynomials [5, 3]. A key new ingredient in the proof is a dimension-independent reverse Hölder inequality for bilinear biforms given in Section 4.3.
Section 3 considers the smallest nontrivial case, that is, the case of cross-positive maps $A:M_{3}(\mathbb{R})\to M_{3}(\mathbb{R})$. We give real algebraic geometry inspired certificates (Nichtnegativstellensätze) for $A$ to be cross-positive; see Theorem 3.4 for the case when $A$ satisfies some mild nonsingularity-type assumption, and Corollary 3.11 for the singular case.
Finally, in Section 5, as a side product of our analysis we provide a randomized polynomial-time algorithm based on semidefinite programming [43] for producing proper cross-positive maps.
2. Preliminaries
2.1. Cross-positivity in the language of operator semigroups
Consider a $\ast$-linear map $A:M_{n}(\mathbb{R})\to M_{n}(\mathbb{R})$. For each $t\in\mathbb{R}$ the linear map $\exp(tA):M_{n}(\mathbb{R})\to M_{n}(\mathbb{R})$ is defined by $\exp(tA)=\sum_{i=0}^{\infty}\frac{1}{i!}(tA)^{i}$. The operator valued function $t\mapsto\exp(tA)$ is the solution of the differential equation $\dot{X}(t)=AX(t)$, which makes it important in analysis and applications to physics [25] and math finance [13, 14]. The well-known formula $\exp((s+t)A)=\exp(sA)\circ\exp(tA)$ implies that the set
$\{\exp(tA)\colon t\geq 0\}$
is a (one-parameter) semigroup under composition. The $\ast$-linear map $A$ is the generator of this one-parameter semigroup. We call $A$ exponentially-positive, resp. completely exponentially-positive, if $\exp(tA)$ is a positive, resp. completely positive map for all $t\geq 0$. In such a case the semigroup $\{\exp(tA)\colon t\geq 0\}$ is a positive, resp. completely positive one-parameter semigroup. Note the positivity of linear maps and their one-parameter semigroups is studied more generally over ordered vector spaces, in finite and infinite dimensions, and for bounded and unbounded linear operators. We refer the reader to [17, 30] for detailed studies.
The (complete) exponential positivity property can be rephrased in a more traditional matrix theory using (complete) cross-positivity.
Theorem 2.1 ([38, Theorem 3]).
A $\ast$-linear map $A:M_{n}(\mathbb{R})\to M_{n}(\mathbb{R})$ is exponentially-positive if and only if it is cross-positive.
Corollary 2.2.
A $\ast$-linear map $A:M_{n}(\mathbb{R})\to M_{n}(\mathbb{R})$ is completely exponentially-positive if and only if it is completely cross-positive.
Proof.
@@ 2.1. Cross-positivity in the language of operator semigroups
By definition, $A$ is completely cross-positive if and only if $I_{k}\otimes A$ is cross-positive for each $k\in\mathbb{N}$. By Theorem 2.1 this holds if and only if $I_{k}\otimes A$ is exponentially-positive for each $k\in\mathbb{N}$, i.e., if and only if
$\forall k\in\mathbb{N},\forall t\geq 0,\forall X\succeq 0$ $:$ $\exp(t(I_{k}\otimes A))(X)=\sum_{i=0}^{\infty}\frac{1}{i!}t^{i}(I_{k}\otimes A)^{i}(X)$
@@
However, this is equivalent to complete positivity of $\exp(tA)$ for each $t\geq 0$, i.e., to complete exponential positivity of $A$. ∎
2.2. Cross-positive maps and biquadratic biforms
Let $n\geq 2$ and let $\mathbb{S}_{n}(\mathbb{R})$ stand for the set of all real symmetric $n\times n$ matrices. To each linear map $A:\mathbb{S}_{n}(\mathbb{R})\to\mathbb{S}_{n}(\mathbb{R})$ we assign the biquadratic form $p_{A}\in\mathbb{R}[{\tt x},{\tt y}]$ as in (1.3). Let $I\subseteq\mathbb{R}[{\tt x},{\tt y}]$ be the ideal generated by ${\tt y}^{T}{\tt x}=\sum_{i=1}^{n}{\tt x}_{i}{\tt y}_{i}$, and let $V(I)$ be the corresponding real variety
@@ 2.2. Cross-positive maps and biquadratic biforms
The variety $V(I)$ is an irreducible hypersurface for $n\geq 2$ and the defining polynomial ${\tt y}^{T}{\tt x}$ changes sign on $\mathbb{R}^{2n}$. Hence the ideal $I$ is real radical [8, Theorem 4.5.1]. Thus $I$ is the vanishing ideal of $V(I)$, i.e., a polynomial $p\in\mathbb{R}[{\tt x},{\tt y}]$ vanishes on $V(I)$ if and only if $p\in I$.
A sum of a positive map and a map of the form
(2.1) $A(X)=CX+XC^{T}\quad\text{for some }C\in M_{n}(\mathbb{R})\text{ and for all }X\in M_{n}(\mathbb{R})$
is clearly cross-positive. The converse is true up to closure, see [38, Lemma 6 and Theorem 2]. It was long conjectured that each cross-positive map is a sum of a positive map and a map of the form (2.1) (see [15, Section 4] or [13, p.409]), but a counterexample was found in [23]. Such counterexamples were called exotic cross-positive maps in [23]. On the other hand, an analogous counterexample does not exist for completely cross-positive maps (see [25, Theorem 3]).
The following is a special case of [23, Corollary 15] and [39, Theorem 2], but can also be established by a straight-forward calculation.
@@
For a linear map $A:\mathbb{S}_{n}(\mathbb{R})\to\mathbb{S}_{n}(\mathbb{R})$ we have $p_{A}\in I$ if and only if it is of the form (2.1) for every $X\in\mathbb{S}_{n}(\mathbb{R})$.
@@
[…]
@@
Let $A:\mathbb{S}_{n}(\mathbb{R})\to\mathbb{S}_{n}(\mathbb{R})$ be a linear map. The following are equivalent:
@@
    AA extends to some completely cross-positive map $\widetilde{A}:M_{n}(\mathbb{R})\to M_{n}(\mathbb{R})$;
@@
Let ${\underline{r}}:=(r_{1},\ldots,r_{n})\in\mathbb{Z}^{n}_{+}$ stand for a tuple of nonnegative integers, ${\tt x}^{\underline{r}}$ for the monomial ${\tt x}_{1}^{r_{1}}\cdots{\tt x}_{n}^{r_{n}}$ and $\operatorname{conv}(E)\subseteq\mathbb{R}^{n}$ for the convex hull of the set $E\subseteq\mathbb{R}^{n}$. Recall that the Newton polytope $N(p)$ of a polynomial $p({\tt x})=\sum_{{\underline{r}}}c_{{\underline{r}}}{\tt x}^{{\underline{r}}}$, where $c_{\underline{r}}\in\mathbb{R}\setminus\{0\}$, is the convex hull of the exponent vectors of the monomials appearing nontrivially in $p$, i.e.,
@@
[…]
@@
where $i,\ell=2,\ldots,n-1$, $j,k=1,\ldots,n$. By [34, Theorem 1], we have the inclusions $N(q^{(\ell)})\subseteq\frac{1}{2}N(q_{A})$ of Newton polytopes, which implies that each $q^{(\ell)}$ is a linear combination of the monomials
@@
Proof of Claim. We consider how each of the monomials in (2.6) can appear in $q^{(\ell)}({\tt x},\alpha)$. For $j=1,\ldots,n$, the monomials
@@
If any of the monomials from (2.7) occurs in $q^{(\ell)}({\tt x},\alpha)$, then it also occurs in the polynomial $q^{(\ell)}(\widehat{\tt x}_{i-1},\widehat{\alpha}_{i-1})$ with the same coefficient. By definition,
@@
Hence, each $q^{(\ell)}(\widehat{{\tt x}}_{i-1},\widehat{\alpha}_{i-1})$ vanishes on $V(\alpha_{i}{\tt x}_{i+1}-\alpha_{i-1}{\tt x}_{i-1})$. Since $\alpha_{i}{\tt x}_{i+1}-\alpha_{i-1}{\tt x}_{i-1}$ is irreducible in $\mathbb{R}[{\tt x},\alpha]$ and it changes sign on $\mathbb{R}^{2n-1}$, it follows by [8, Theorem 4.5.1] that
@@
where $p_{i-1}$ is a linear form in $\widehat{{\tt x}}_{i-1}$. Now (2.8) implies that the monomials from (2.7) can appear nontrivially in $q^{(\ell)}(\widehat{{\tt x}}_{i-1},\widehat{\alpha}_{i-1})$ only from the scalar multiple of the term
@@
Using the Claim and (2.5) it follows that $p_{A}({\tt x},{\tt y})$ agrees on a dense subset of $V(I)$ and by continuity on the whole $V(I)$ with a sum of squares polynomial, which we denote by $r({\tt x},{\tt y})$. Since $p_{A}-r$ vanishes on $V(I)$, the polynomial ${\tt y}^{T}{\tt x}$ is irreducible in $\mathbb{R}[{\tt x},{\tt y}]$ and its sign changes on $\mathbb{R}^{2n-1}$, it follows that $p_{A}-r\in I$ by [8, Theorem 4.5.1]. ∎